OpenStudy (anonymous):

|3x-5| less than or equal to 8

OpenStudy (lxelle):

Let 3x-5 = 8 or 3x-5 = -8

OpenStudy (lxelle):

So, 3x=13 or 3x= -3 x= 13/3 or x= -1 therefore -1<x<13/3